OpenStudy (anonymous):

Solve: -2(2x + 5) - 3 = -3(x - 1) I don't understand how to do this.....

OpenStudy (anonymous):

-16 -13 -10 16

OpenStudy (anonymous):

Those are the answers it gives me

hero (hero):

Hold on a second. I can find out for sure what the answer is

OpenStudy (anonymous):

Okay

hero (hero):

-16

OpenStudy (anonymous):

How did you get that?

hero (hero):

Hold on, I see the mistake I made. Let me show you:

OpenStudy (anonymous):

okay

hero (hero):

-2(2x + 5) - 3 = -3(x - 1) -4x-10-3 = -3x+3 -4x-13=-3x+3 -13-3 = 4x-3x -16 = x

hero (hero):

What level of math are you taking?

OpenStudy (anonymous):

Algebra 1

hero (hero):

Are you looking for a personal tutor?

OpenStudy (anonymous):

Kinda, i don't really know how to do this math. And i dont have a teacher cuz im homechooled

hero (hero):

That's what I figured. I can help you :D

OpenStudy (anonymous):

Okie dokie (: I mean i've already failed algebra 1 already :/

hero (hero):

I understand. No worries. I can help you get up to speed. Do you have skype?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Okay i did but my cam is broken